OpenStudy (anonymous):

find all solutions of 2-2cos^2x=sinx+1

OpenStudy (campbell_st):

well make the substitution for cos^2 = 1 - sin^2 so you have 2 - 2(1 - sin^2(x)) = sin(x) + 1 which becomes after simplifying and rewriting 2sin^2(x) - sin(x) - 1 = 0 which is a quadratic that can be factored. when you have the factors, you can solve for x.